OpenStudy (anonymous):

please help; what is the equation of a line with the given points are (4,-1), (-2,-6)

OpenStudy (amorfide):

to get the equation of a line you do y-y1=m(x-x1) y1 is a y coordinate x1 is an x coordinate

OpenStudy (amorfide):

to get the gradient 

OpenStudy (amorfide):

y2 and y1 being the y coordinates x2 and x1 being the x coordinates

OpenStudy (anonymous):

find slope first (y1-y2) / (x1-x2) so it is (-6 - (-1)) / (-2 - 4) = 5/6 y = mx + b y= 5/6x + b -1 = 5/6(4) + b solvve for b b = -26/6 eq: y= 5/6x -26/6

OpenStudy (anonymous):

alright i did that and got y=-7/6 x + 11/3

OpenStudy (anonymous):

to jayz657, i think its y2-y1/x2-x1..... to find slope right?

OpenStudy (anonymous):

doesnt matter you get slope either way y1 and y2 is to show these are 2 different numbers at 2 different points

OpenStudy (anonymous):

aight

OpenStudy (anonymous):

shount it be -5/6

OpenStudy (anonymous):

its positive for me check your signs

OpenStudy (anonymous):